OpenStudy (anonymous):

p(x)=-2x^2+60x-120 Im not sure how to do this im getting so lost on my homework any help?

OpenStudy (saifoo.khan):

what's the question?

OpenStudy (anonymous):

Im not really sure where to start I need to find 3 things Whats the largest number of cases that can be sold? How its possible to loose money? how many cases should be sold to make a profit? Im not sure where to start in finding these three options.

OpenStudy (anonymous):

\[\text{saifoo}_{100}\] to the rescue

OpenStudy (anonymous):

Divide both sides by -2: x^2-30 x+60 = 0 Subtract 60 from both sides: x^2-30 x = -60 Add 225 to both sides: x^2-30 x+225 = 165 Factor the left hand side: (x-15)^2 = 165 Take the square root of both sides: sqrt(x-15) = sqrt(165) Eliminate the absolute value: x-15 = -sqrt(165) or x-15 = sqrt(165) Add 15 to both sides: x = 15-sqrt(165) or x-15 = sqrt(165) Add 15 to both sides: x = 15-sqrt(165) or x = 15+sqrt(165)

OpenStudy (anonymous):

i don't think it asked for the zeros just the max use \[x=-\frac{b}{2a}=-\frac{60}{2\times -2}=\frac{60}{4}=15\]

OpenStudy (anonymous):

hmmm where did the 225 come from?

OpenStudy (saifoo.khan):

is the question compelte?

OpenStudy (saifoo.khan):

complete*

OpenStudy (anonymous):

that's the whole equation in my book unfortunatly i'm taking an online stats class i'm lost and my teacher is no help.

OpenStudy (saifoo.khan):

there must be something more a paragraph or something..

OpenStudy (anonymous):

Pat Kan owns a factory that manufactures souvenir key chains. Her weekly profit ( in hundreds of dollars) is given by p(x)=-2x^2+60x-120, where x is the number of cases of key chains sold.

hero (hero):

Whats the largest number of cases that can be sold? -b/2a How its possible to loose money? when x generates a P(x) > 0 how many cases should be sold to make a profit? When x generates a P(x) > 0

hero (hero):

You can find the exact number of cases through trial and error. Start by graphing the function

hero (hero):

Or you can do what Dhashni did

OpenStudy (anonymous):

yours seems a bit more simplistic i'm looking at what my teacher has for us and all its doing is making me more lost ... I think i get it im gonna try it thanks for the help to all

hero (hero):

And also, If you want to factor -2x^2-60x-120: -2(x^2-30x-60) -2(x^2-30x-30x-60) -2(x(x-30)-30(x-30) -2((x-30)(x-30)) -2(x-30)^2

hero (hero):

Oh and I made a mistake. For loss, x has to generate a P(x) < 0